# QiMeng-CRUX: Narrowing the Gap between Natural Language and Verilog via Core Refined Understanding eXpression

### v2

@@ Conclusion
In this work, we present CRUX, a structured intermediate space designed to bridge the gap between free-form natural language descriptions and formal Verilog code. By explicitly modeling Module Interfaces, Core Functions, and Key Considerations, CRUX captures essential design intent and provides clear, targeted guidance for downstream code generation. Through a two-stage framework combining Joint Expression Modeling and Dual-Space Optimization, our method not only improves generation accuracy but also enhances the robustness and interpretability of large language models in hardware design tasks. Extensive evaluations show that QiMeng-CRUX-V achieves state-of-the-art performance across multiple Verilog benchmarks, demonstrating the effectiveness of semantically structured guidance in enabling precise and synthesizable HDL generation.
@@ Appendix A Appendix
1) Removing KL Divergence. QiMeng-CRUX-V constructs an intermediate space that bridges free-form natural language descriptions and domain-specific Verilog code. To ensure the effectiveness of this space for precise Verilog generation, we introduce a CRUX-Reward, which directly optimizes the semantic alignment between the CRUX and correct Verilog codes. Consequently, we do not employ KL-divergence to regularize the policy model. Removing the KL constraint encourages the model to generate more diverse CRUXes and Verilog codes, while CRUX-Reward and Code-Reward respectively serve as the primary objectives that guide the generation toward correctness and faithfulness, providing a more direct and effective optimization mechanism.
To further validate the effectiveness of this design, we conduct an ablation study based on QiMeng-CRUX-V-SFT, examining the impact of removing the KL-divergence term. The evaluation is performed on three benchmarks: VerilogEval-V1-Human, VerilogEval-V2-SR, and RTLLM-v2, and we report pass@1 under identical training steps and datasets. The results are summarized as follows.
@@
Table 4: Ablation study on the KL-divergence penalty in Stage II of QiMeng-CRUX-V training across three benchmarks, showing consistent performance gains when the KL penalty is removed.
@@
During GRPO training, we save a model checkpoint every 10 steps. Figure7 illustrates the evolution of the pass@k metric on the Spec-to-RTL benchmark from VerilogEval-V2 as training progresses. As shown, pass@1 steadily increases over time but experiences a slight drop after around 460 steps, while pass@10 decreases initially and then shows a mild improvement after 400 steps. The red star marks the best performance among the steps, and the green dashed line indicates the final model we selected. We adopt the checkpoint at 420 steps as our QiMeng-CRUX-V-Final model—not the best on Spec-to-RTL alone, but demonstrating the most balanced performance across multiple benchmarks.
@@
In the main results, we demonstrate that the CRUX space produced by QiMeng-CRUX-V-Final on the Spec-to-RTL benchmark can serve either as a complete prompt or as a bridging component to the original design prompt. This shows that the learned CRUX space provides transferable benefits to general-purpose code models, without requiring any additional training. To further support this observation, we evaluate the transferability of CRUX on two additional benchmarks that also reflect real-world hardware design scenarios: VerilogEval-V1-Human, which features hand-crafted and semantically diverse prompts, and RTLLM-V2, which involves high-level and complex specifications resembling practical workflows. Full results are presented in Table 6.

### v3

@@ Conclusion
In this work, we present CRUX, a structured intermediate space designed to bridge the gap between free-form natural language descriptions and formal Verilog code. By explicitly modeling Module Interfaces, Core Functions, and Key Considerations, CRUX captures essential design intent and provides clear, targeted guidance for downstream code generation. Through a two-stage framework combining Joint Expression Modeling and Dual-Space Optimization, our method not only improves generation accuracy but also enhances the robustness and interpretability of large language models in hardware design tasks. Extensive evaluations show that QiMeng-CRUX achieves state-of-the-art performance across multiple Verilog benchmarks, demonstrating the effectiveness of semantically structured guidance in enabling precise and synthesizable HDL generation.
@@ Appendix A Appendix
1) Removing KL Divergence. QiMeng-CRUX constructs an intermediate space that bridges free-form natural language descriptions and domain-specific Verilog code. To ensure the effectiveness of this space for precise Verilog generation, we introduce a CRUX-Reward, which directly optimizes the semantic alignment between the CRUX and correct Verilog codes. Consequently, we do not employ KL-divergence to regularize the policy model. Removing the KL constraint encourages the model to generate more diverse CRUXes and Verilog codes, while CRUX-Reward and Code-Reward respectively serve as the primary objectives that guide the generation toward correctness and faithfulness, providing a more direct and effective optimization mechanism.
To further validate the effectiveness of this design, we conduct an ablation study based on QiMeng-CRUX-SFT, examining the impact of removing the KL-divergence term. The evaluation is performed on three benchmarks: VerilogEval-V1-Human, VerilogEval-V2-SR, and RTLLM-v2, and we report pass@1 under identical training steps and datasets. The results are summarized as follows.
@@
Table 4: Ablation study on the KL-divergence penalty in Stage II of QiMeng-CRUX training across three benchmarks, showing consistent performance gains when the KL penalty is removed.
@@
During GRPO training, we save a model checkpoint every 10 steps. Figure7 illustrates the evolution of the pass@k metric on the Spec-to-RTL benchmark from VerilogEval-V2 as training progresses. As shown, pass@1 steadily increases over time but experiences a slight drop after around 460 steps, while pass@10 decreases initially and then shows a mild improvement after 400 steps. The red star marks the best performance among the steps, and the green dashed line indicates the final model we selected. We adopt the checkpoint at 420 steps as our QiMeng-CRUX-Final model—not the best on Spec-to-RTL alone, but demonstrating the most balanced performance across multiple benchmarks.
@@
In the main results, we demonstrate that the CRUX space produced by QiMeng-CRUX-Final on the Spec-to-RTL benchmark can serve either as a complete prompt or as a bridging component to the original design prompt. This shows that the learned CRUX space provides transferable benefits to general-purpose code models, without requiring any additional training. To further support this observation, we evaluate the transferability of CRUX on two additional benchmarks that also reflect real-world hardware design scenarios: VerilogEval-V1-Human, which features hand-crafted and semantically diverse prompts, and RTLLM-V2, which involves high-level and complex specifications resembling practical workflows. Full results are presented in Table 6.